# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 3.2 The role of closed channels
In principle, the expansion (8) may include internal states with excitation energies below ($\epsilon_{i}<E$) or above ($\epsilon_{i}>E$) the total energy, which are referred to as open and closed channels, respectively. Closed channels violate energy conservation and, as such, cannot contribute asympotically to the scattering wavefunction. However, due to the couplings with other channels, they can affect the open channels and hence the cross sections.
The solution of the CDCC equations (13) in the presence of closed channels is totally analogous to the case with open channels, except for the modification of the boundary conditions of the radial functions associated with closed channels, which now reads
@@
where $W_{-\eta_{i},L+1/2}$ is a Whittaker function with $\eta$ the Sommerfeld parameter, defined as $\eta=Z_{p}Z_{t}e^{2}/\hbar v$. In general, the effect of closed channels is small at sufficiently high energies. However, at low energies, they can have a strong influence on the scattering observables. An example is shown in Fig. 6, corresponding to the reaction $d+^{12}\mathrm{C}\rightarrow p+n+^{12}\mathrm{C}$ at 12 MeV Oga16 . CDCC calculations with and without the inclusion of closed channels are compared with Faddeev calculations, which account also for the effect of closed channels. The upper and lower panels are for the angular and excitation energy distributions. The importance of closed channels is clearly seen.
Figure 6: (a) Angular distribution and (b) breakup energy distribution of the elastic breakup cross section for 12C($d,pn$)12C at 12 MeV. The solid, dashed, and dash-dotted lines in each panel show the converged CDCC result, the result of the CDCC method calculated with only the open channels, and the result of FAGS theory taken from Ref. Upa12 , respectively. The dotted line in (a) is the same as the solid line but omitting the odd partial waves between $p$ and $n$. Taken from Oga16 , with authorization from APS.
@@ 3.3 Inclusion of core excitations
Excitations of the projectile constituents ($b$ and $x$ in our case) may take place along with the projectile breakup. This mechanism is neglected in the standard formulation of the CDCC method. For example, for the scattering of halo nuclei, the core fragment $b$ is assumed to be inert and, as such, the projectile states are described by pure single-particle or pure cluster states. Except for very specific cases (e.g. when $b$ is an alpha particle), the structure of a composite system of the form ($a=b+x$) is known to contain in general admixtures of core-excited components. Additionally, the interaction of the core with the target will produce excitations and deexcitations of the former during the collision, thereby modifying the reaction observables to some extent. These two effects (structural and dynamical) have been investigated within extended versions of the DWBA and CDCC methods Cre11 ; Mor12 ; Sum06 ; Die14 . For example, considering only possible excitations of $b$, the effective three-body Hamiltonian is generalized as follows:
@@
where $n$ is an index labeling the states with angular momentum $J$ (with projection $M$), $\alpha\equiv\{\ell,s,j,I\}$, with $I$ the core intrinsic spin, $\vec{j}=\vec{\ell}+\vec{s}$ and $\vec{j}_{p}=\vec{j}+\vec{I}$. The functions $\Phi_{I}(\xi_{b})$ and $\varphi_{\alpha}(\vec{r})$ describe, respectively, the core states and the valence–core relative motion. This expansion is in principle also valid for continuum states, in which case a procedure of continuum discretization must be used.
Once the projectile states (21) have been calculated, the three-body wave function is expanded in a basis of such states, as in the standard CDCC method (c.f. Eq. (12)),
@@
Early calculations using this extended CDCC method (referred to as XCDCC) were first performed by Summers et al. Sum06 ; Sum07 for 11Be and 17C on 9Be and 11Be+$p$, finding a very little core excitation effect in all these cases. However, later calculations for the 11Be+$p$ reaction based on a alternative implementation of the XCDCC method using a pseudo-state representation of the projectile states Die14 suggested much larger effects. The discrepancy was found to be due to an inconsistency in the numerical implementation of the XCDCC formalism presented in Ref. Sum06 , as clarified in Sum14 . For heavier targets, such as 64Zn or 208Pb, the calculations of Die14 suggest that the core excitation mechanism plays a minor role, although its effect on the structure of the projectile is still important.
Figure 7: Breakup angular distributions for the relative-energy cuts indicated in the labels. Adapted from Ref. Die14 ; Die17 .
@@
The importance of the deformation on the structure of the projectile is clearly evidenced in the elastic and inelastic scattering of 11Be on 197Au at near-barrier energies Pes17 , as shown in Fig. 8. XCDCC calculations based on the particle-plus-rotor model of Ref. Tar03 (solid lines) are able to reproduce simultaneously the elastic, inelastic and breakup angular distributions. By contrast, standard CDCC calculations using single-particle wave functions fail to describe the elastic and inelastic data, despite describing well the breakup. This is due to the overestimation of the $B(E1)$ connecting the ground state with the bound excited state, as discussed in Pes17 .
@@ 3.4 Inclusion of target excitations
In addition to the excitations of the projectile constituents, excitations of the target nucleus may also take place and compete with the projectile breakup mechanism. Note that, within CDCC, the projectile breakup is treated as a inelastic excitation of the projectile to its continuum states and, thus, inclusion of target excitation amounts at including, simultaneously, projectile plus target excitations so their relative importance, and mutual influence, can be assessed. These target excitations can be treated with the collective models mentioned in the previous section. It is worth noting that, within this three-body reaction model, target excitation arises from the non-central part of the valence-target and core-target interactions. To incorporate this effect, the effective Hamiltonian, eq. (5), must be now generalized as:
@@
Recently, the problem has been also addressed by some authors Chau11 ; Gom17a , also in the context of deuteron elastic and inelastic scattering. A recent application of the formalism is shown in Fig. 10, which corresponds to the reaction ${}^{24}{\rm Mg}(d,d)^{24}{\rm Mg}^{*}$ at $E_{d}=70$ MeV, including the ground and first excited states of ${}^{24}{\rm Mg}$, in addition to the deuteron breakup. The data are from Ref. Kis76 . The target excitation was treated within the collective model, using a quadrupole deformation parameter of $\beta_{2}=0.5$. Also included are Faddeev calculations performed by A. Deltuva Del16 . Both calculations reproduce equally well the elastic differential cross section. The calculated inelastic angular distributions are slightly out of phase with the data, but they agree well with each other, pointing to some inadequacy of the structure or potential inputs. Although these inelastic cross section can be also well reproduced with standard DWBA calculations based on a deformed deuteron-target potential, it was shown in Del16 that the extracted deformation parameter obtained with the three-body approach is more consistent with that derived from nucleon-nucleus inelastic scattering.
Figure 10: Elastic (upper) and $(2^{+})$ excitation (lower) differential angular cross sections for $d$+24Mg $E_{d}=70$ ${\rm MeV}$. CDCC calculations are compared with Faddeev calculations from ref. Del16 and with the data of ref. Kis76 . The calculations employ CH89 CH89 parameterization for the $p$+24Mg and $n$+24Mg potentials and treat the target excitation with a collective model, assuming a deformation of $\beta=0.5$ for the 24Mg nucleus. The plot is adapted from ref. Gom17a .
@@ 3.5 Four-body CDCC
Figure 11: Schematic representation of the four-body CDCC, where the projectile is described in Jacobi coordinates $\{\vec{x},\vec{y}\}$. Adapted from Ref. casal15 .
In simple terms, the generalization implies the inclusion of three-body projectile wave functions in the channel states of Eq. (12). These can be written in Jacobi coordinates $\left\{\vec{x},\vec{y}\right\}$ (see Fig. 11), in such a way that the coupled basis (15) becomes
@@
Here, $\eta=\left\{K,l_{x},l_{y},l,S_{x},j,I\right\}$ are the possible sets of quantum numbers associated to the orbital angular momenta and spins coupled to a total $J$, where $K$ is the so-called hypermomentum, and $\varphi_{K}^{l_{x},l_{y}}(\alpha)$ are obtained in terms of Jacobi polynomials. The three-body hyperradial functions $u_{n\eta}(\rho)$ can be obtained within different discretization methods, including the binning procedure manoli09 ; Cub12 and the pseudostate method with different bases manoli08 ; Desc15 ; casal15 . This is done for a fixed value of $K_{max}$, up to convergence. Note that the above wave-function expansion and its coupling order is not unique, and other choices exist in the literature.
Figure 12 shows an application of the four-body CDCC method to describe the ${}^{11}\text{Li}+{{}^{208}}\text{Pb}$ reaction at near-barrier energies Cub12 , where the projectile states and the corresponding coupling potentials are computed with a ${}^{9}\text{Li}+n+n$ model using the binning discretization method. In that work, it is shown that the coupling to low-energy continuum states is crucial to explain the rapid decrease of the elastic differential cross section, and the calculations indicate the presence of a low-lying dipole resonance in 11Li for a proper description of the experimental data.
Figure 12: Experimental ${}^{11}\text{Li}+{{}^{208}}\text{Pb}$ elastic cross section at $E_{lab}=29.8$ MeV, relative to Rutherford, compared to four-body CDCC calculations (solid) and the result without the coupling to continuum states (dashed). Adapted from Ref. Cub12 .
Figure 13 presents results for the ${}^{11}\text{Li}(p,p^{\prime})$ reaction, focusing on the breakup cross section extracted from the four-body CDCC Matsumoto19 . These calculations employ pseudostates to describe the 11Li states, and the so-called complex scaling method is adopted to build the corresponding energy distributons. Again, a dipole resonance is found to be crucial for a proper understanding of the data. Other calculations for the same reaction exist in the literature Descouvemont20 .

### v2

@@ 3.2 The role of closed channels
In principle, expansion (9) may include internal states with excitation energies below ($\varepsilon_{i}<E$) or above ($\varepsilon_{i}>E$) the total energy, which are referred to as open and closed channels, respectively. Closed channels violate energy conservation and, as such, cannot contribute asymptotically to the scattering wavefunction. However, due to the couplings with other channels, they can affect the open channels and hence the cross sections.
The solution of the CDCC equations (14) in the presence of closed channels is totally analogous to the case with open channels, except for the modification of the boundary condition of the radial functions associated with closed channels, which now reads
@@
where $W_{-\eta_{i},L+1/2}$ is a Whittaker function with $\eta$ the Sommerfeld parameter, defined as $\eta=Z_{p}Z_{t}e^{2}/\hbar v$. In general, the effect of closed channels is small at sufficiently high energies. However, at low energies, they can have a strong influence on the scattering observables. An example is shown in Fig. 7, corresponding to the reaction $d+^{12}\mathrm{C}\rightarrow p+n+^{12}\mathrm{C}$ at 12 MeV Oga16 . CDCC calculations with and without the inclusion of closed channels are compared with Faddeev calculations performed in the momentum-space formulation of Alt, Grassberger and Sandhas (FAGS) AGS , which provide an essentially exact solution of the same Hamiltonian and therefore also account for the effect of closed channels. The upper and lower panels are for the angular and excitation energy distributions. The importance of closed channels is clearly seen. Only when closed channels are included, the CDCC calculation approaches the Faddeev solution.
Figure 7: (a) Angular distribution and (b) breakup energy distribution of the elastic breakup cross section for 12C($d,pn$)12C at 12 MeV. The solid, dashed, and dash-dotted lines in each panel show the converged CDCC result, the result of the CDCC method calculated with only the open channels, and the result of the Faddeev-AGS (FAGS) theory taken from Ref. Upa12 , respectively. The dotted line in (a) is the same as the solid line but omitting the odd partial waves between $p$ and $n$. Taken from Oga16 , with authorization from APS.
3.3 Trivially equivalent polarization potential
From a coupled-channel calculation (in particular, CDCC) one may extract an effective polarization potential, also called trivial equivalent local polarization potential (TELP). For simplicity, we assume that there is only one entrance channel ($\beta_{i}=\beta_{0}$) so that the label $\beta_{i}$ in (20) can be omitted. Then the radial equation for the elastic channel becomes:
$\left(-\frac{\hbar^{2}}{2\mu}{d^{2}\over dR^{2}}+\frac{\hbar^{2}L_{0}(L_{0}+1)}{2\mu R^{2}}+U_{\beta_{0},\beta_{0}}^{J_{T}}(R)+\varepsilon_{0}-E\right)\chi^{J_{T}}_{\beta_{0}}(R)$
$=-\sum_{\beta^{\prime}}U_{\beta_{0},\beta^{\prime}}^{J_{T}}(R)\chi^{J_{T}}_{\beta^{\prime}}(R)$ $=0$ (20)
The right-hand side of this equation is then used to define the angular-momentum dependent polarization potential:
$U_{\text{pol}}(R)=\frac{\sum_{J_{T}}w_{J_{T}}(R)U_{J_{T}}^{\text{TELP}}(R)}{\sum_{J_{T}}w_{J_{T}}(R)},$ (21)
where $U_{J_{T}}^{TELP}(R)$ is the “trivial equivalent potential” for the total angular momentum $J_{T}$ defined by
$U_{J_{T}}^{TELP}(R)=\frac{1}{\chi^{J_{T}}_{\beta_{0}}(R)}\sum_{\beta^{\prime}\neq\beta_{0}}U_{\beta_{0},\beta^{\prime}}^{J_{T}}(R)\chi^{J_{T}}_{\beta^{\prime}}(R),$ (22)
and $w_{J_{T},\beta_{i}}(R)$ are weight factors chosen as
$w_{J_{T}}(R)=(2J_{T}+1)(1-\mid S^{J_{T}}_{\beta_{0},\beta_{0}}\mid^{2})\mid\chi^{J_{T}}_{\beta_{0}}(R)\mid^{2},$ (23)
where $S^{J_{T}}_{\beta_{0},\beta_{0}}$ are the elastic S-matrix elements for each $J_{T}$ value. A single channel calculation (elastic channel) using the sum of “$U_{\beta_{0},\beta_{0}}^{J_{T}}(R)$ + $U_{\text{pol}}(R)$” should approximately reproduce the elastic scattering cross sections.
In Fig. 9, we show for illustration the polarization potential derived from the CDCC calculation for the reaction d+58Ni reaction at 80 MeV (c.f. Fig. 4). For comparison, we show the ground state diagonal monopole coupling potential ($U_{00}$) In this case, the effect the coupling to the breakup channels produces is mostly repulsive and absorptive.
Figure 8: Average polarization potential $U_{\text{pol}}$ (solid lines), compared with the ground state diagonal monopole potential $U_{00}$ (dashed) for the d+58Ni reaction at 80 MeV.
@@ 3.4 Inclusion of core excitations
Excitations of the projectile constituents ($b$ and $x$ in our case) may take place along with projectile breakup. This mechanism is neglected in the standard formulation of the CDCC method. For example, for the scattering of halo nuclei, the core fragment $b$ is assumed to be inert and, as such, the projectile states are described by pure single-particle or pure cluster states. Except for very specific cases (e.g. when $b$ is an alpha particle), the structure of a composite system of the form ($a=b+x$) is known to contain in general admixtures of core-excited components. Additionally, the interaction of the core with the target will produce excitations and deexcitations of the former during the collision, thereby modifying the reaction observables to some extent. These two effects (structural and dynamical) have been investigated within extended versions of the DWBA and CDCC methods Cre11 ; Mor12 ; Sum06 ; Die14 . For example, considering only possible excitations of $b$, the effective three-body Hamiltonian is generalized as follows:
@@
where $n$ is an index labeling the states with angular momentum $J$ (with projection $M$), $\alpha\equiv\{\ell,s,j,I\}$, with $I$ the core intrinsic spin, $\vec{j}=\vec{\ell}+\vec{s}$ and $\vec{J}=\vec{j}+\vec{I}$. The functions $\Phi_{I}(\xi_{b})$ and $\varphi_{\alpha}(\vec{r})$ describe, respectively, the core states and the valence–core relative motion. This expansion is in principle also valid for continuum states, in which case a procedure of continuum discretization must be used.
Once the projectile states in Eq. (26) have been calculated, the three-body wave function is expanded in a basis of such states, as in the standard CDCC method (c.f. Eq. (13)),
@@
The first calculations using this extended CDCC method (referred to as XCDCC) were performed by Summers et al. Sum06 ; Sum07 for 11Be and 17C on 9Be and 11Be+$p$, finding a very little core excitation effect in all these cases. However, later calculations for the 11Be+$p$ reaction based on a alternative implementation of the XCDCC method using a pseudo-state representation of the projectile states Die14 suggested much larger effects. The discrepancy was found to be due to an inconsistency in the numerical implementation of the XCDCC formalism presented in Ref. Sum06 , as clarified in Sum14 . For heavier targets, such as 64Zn or 208Pb, the calculations of Die14 suggest that the core excitation mechanism plays a minor role in the reaction dynamics, although its effect on the structure of the projectile is still important.
Figure 9: Breakup angular distributions for the relative-energy cuts indicated in the labels. Dashed line are the calculations in which the 10Be+p potential includes only a central part and hence ignores possible excitations of 10Be. Solid lines are the full XCDCC calculations, including 10Be excitations within a collective model. Adapted from Ref. Die14 ; Die17 .
@@
The importance of the deformation on the structure of the projectile is clearly evidenced in the elastic and inelastic scattering of 11Be on 197Au at near-barrier energies Pes17 , as shown in Fig. 10. XCDCC calculations based on the particle-plus-rotor model of Ref. Tar03 (solid lines) are able to reproduce simultaneously the elastic, inelastic and breakup angular distributions. In contrast, standard CDCC calculations using single-particle wave functions fail to describe the elastic and inelastic data, despite describing well the breakup. This is due to the overestimation of the $B(E1)$ connecting the ground state with the bound excited state, as discussed in Pes17 .
@@ 3.5 Inclusion of target excitations
In addition to the excitations of the projectile constituents, excitations of the target nucleus may also take place and compete with the projectile breakup mechanism. Note that, within CDCC, projectile breakup is treated as a inelastic excitation of the projectile to its continuum states and, thus, inclusion of target excitation amounts to including, simultaneously, projectile plus target excitations so their relative importance, and mutual influence, can be assessed. These target excitations can be treated with the collective models mentioned in the previous section. It is worth noting that, within this three-body reaction model, target excitation arises from the non-central part of the valence-target and core-target interactions. To incorporate this effect, the effective Hamiltonian, Eq. (6), must be now generalized as Yah86 :
@@
Recently, the problem has also been addressed by some authors Chau11 ; Gom17a , also in the context of elastic and inelastic deuteron scattering. A recent application of the formalism is shown in Fig. 12, which corresponds to the reaction ${}^{24}{\rm Mg}(d,d)^{24}{\rm Mg}^{*}$ at $E_{d}=70$ MeV, including the ground and first excited states of ${}^{24}{\rm Mg}$, in addition to the deuteron breakup. The data are from Ref. Kis76 . The target excitation was treated within the collective model, using a quadrupole deformation parameter of $\beta_{2}=0.5$. Also included are Faddeev calculations performed by A. Deltuva Del16 . Both calculations reproduce equally well the elastic differential cross section. The calculated inelastic angular distributions are slightly out of phase with the data, but they agree well with each other, pointing to some inadequacy of the structure or potential inputs.
Note that the traditional procedure to analyze these reactions employs DWBA or coupled channel methods, based on a deformed deuteron-target potential. However, it is not guaranteed that the deformation parameters extracted by this procedure are consistent with those derived from nucleon-nucleus inelastic scattering. The three-body approaches outlined in this section (CDCC and Faddeev) have the advantage of treating both types of reactions (deuteron-nucleus and nucleon-nucleus) from the same underlying (deformed) nucleon-nucleus potentials. In fact, it was shown in Del16 that the extracted deformation parameter obtained with the three-body approach is more consistent with that derived from nucleon-nucleus inelastic scattering.
Figure 12: Elastic (upper) and $(2^{+})$ excitation (lower) differential angular cross sections for $d$+24Mg at $E_{d}=70$ ${\rm MeV}$. CDCC calculations are compared with Faddeev calculations from Ref. Del16 and with the data of Ref. Kis76 . The calculations employ CH89 CH89 parameterization for the $p$+24Mg and $n$+24Mg potentials and treat the target excitation with a collective model, assuming a deformation of $\beta=0.5$ for the 24Mg nucleus. The plot is adapted from Ref. Gom17a .
@@ 3.6 Four-body CDCC
Figure 13: Schematic representation of the four-body CDCC coordinate system, where the projectile is described in Jacobi coordinates $\{\vec{x},\vec{y}\}$. Adapted from Ref. casal15 .
In simple terms, the generalization implies the inclusion of three-body projectile wave functions in the channel states of Eq. (13). These can be written in Jacobi coordinates $\left\{\vec{x},\vec{y}\right\}$ (see Fig. 13), in such a way that the coupled basis in Eq. (16) becomes
@@
Here, $\eta=\left\{K,l_{x},l_{y},l,S_{x},j,I\right\}$ are the possible sets of quantum numbers associated with the orbital angular momenta and spins coupled to a total $J$, where $K$ is the so-called hypermomentum, and $\varphi_{K}^{l_{x},l_{y}}(\alpha)$ are obtained in terms of Jacobi polynomials. The three-body hyperradial functions $u_{n\eta}(\rho)$ can be obtained within different discretization methods, including the binning procedure manoli09 ; Cub12 and the pseudostate method with different bases manoli08 ; Desc15 ; casal15 . This is done for a fixed value of $K_{max}$, up to convergence. Note that the above wave-function expansion and its coupling order is not unique, and other choices exist in the literature.
Figure 14 shows an application of the four-body CDCC method to the ${}^{6}\text{He}+{{}^{208}}\text{Pb}$ reaction at 22 MeV. The combined data from two experiments San08 ; Aco11 are compared with four-body CDCC calculations (solid line) from Ref. Rod09 assuming a ${}^{4}\text{He}+n+n$ model for the 6He states and using the binning discretization method for the continuum. The dashed line is the CDCC calculation omitting the coupling to the unbound states of the projectile. The comparison of the two curves evidences the importance of the breakup channels on the elastic cross section.
Figure 14: Elastic differential cross section, relative to Rutherford, for the 6He+208Pb reaction at 22 MeV. The data are from Refs. San08 ; Aco11 . The lines are four-body CDCC calculations, using a binning discretization procedure, from Ref. Rod09 . The dashed line is the CDCC calculation omitting the breakup channels.
As discussed in the introductory section (see also Rod08 ) a significant part of the suppression of the elastic scattering cross section stems from the Coulomb dipole polarizability of the projectile which, in the CDCC calculation, is included in the coupling potentials connecting the projectile ground state with the $J=1^{-}$ continuum states. Thus, the effect is expected to be enhanced for more weakly bound halo nuclei, which exhibit a larger $E1$ strength at low excitation energies. This is confirmed in Fig. 15, which shows the experimental and calculated elastic differential cross section for the ${}^{11}\text{Li}+{{}^{208}}\text{Pb}$ reaction at a near-barrier energy Cub12 . In that work, it is shown that the coupling to low-energy continuum states is crucial to explain the rapid decrease of the elastic differential cross section, and the calculations indicate the presence of a low-lying dipole resonance in 11Li for a proper description of the experimental data.
Figure 15: Experimental ${}^{11}\text{Li}+{{}^{208}}\text{Pb}$ elastic cross section at $E_{\text{lab}}=29.8$ MeV, relative to Rutherford, compared to four-body CDCC calculations (solid) and the result without the coupling to continuum states (dashed). Adapted from Ref. Cub12 .
Figure 16 presents results for the ${}^{11}\text{Li}(p,p^{\prime})$ reaction, focusing on the breakup cross section extracted from four-body CDCC Matsumoto19 . These calculations employ pseudostates to describe the 11Li states, and the so-called complex scaling method is adopted to build the corresponding energy distributons. Again, a dipole resonance is found to be crucial for a proper understanding of the data. Other calculations for the same reaction exist in the literature Descouvemont20 . Note that the complex-scaling method has been applied in four-body CDCC calculations for other projectiles, such as ${}^{9}\text{C}({{}^{7}}\text{Be}+p+p)$ Singh21 , thus providing a valuable tool in this kind of studies.
@@
In Sec. 2 it was shown that, in order to reproduce the elastic cross section of halo nuclei in terms of the conventional optical model framework, one requires an unusually large diffuseness parameter for the imaginary part of the potential. This can be linked to the CDCC predictions for these systems making use of the local polarization potential introduced in Sec. 3.3 which is meant to encode the net effect of the full coupling scheme on elastic scattering. As an illustration, in Fig. 18 we compare the phenomenological potential of Table 1 with the effective potential ($U_{\mathrm{00}}+U_{\mathrm{pol}}$) extracted from the four-body CDCC calculations of Rod09 for ${}^{6}\text{He}+{{}^{208}}\text{Pb}$ at $E_{\mathrm{lab}}=22$ MeV. The top and bottom panels correspond to the real and imaginary parts, respectively. The arrows indicate the radii of sensitivity of the real and imaginary parts of the optical potential, according to the OM analysis performed in San08 . It is seen that both the phenomenological and CDCC potentials display similar behavior, and, in particular, the effective potential derived from CDCC has the long-range imaginary contribution obtained in the phenomenological OM analysis.
Figure 18: Bare $U_{00}$ (dotted line) and effective potential, $U_{\mathrm{eff}}=U_{\mathrm{00}}+U_{\mathrm{pol}}$ (dot-dashed line), for 6He+208Pb at 22 MeV, extracted from the CDCC. The dashed line is the phenomenological optical model potential listed in Table 1. The top and bottom panels correspond to the real and imaginary parts, respectively. The vertical arrows denote the radius of maximum sensitivity, as estimated in San08 .